SUMMER 2003 REU PARTICIPANTS AND ASSIGNMENTS

Dean Baskin Lauren Beitler Emily Butler Joyce Chen Zachary Cohn Nathan Czuba Keith D’Souza Matthew Gelvin Joanna Goldenberg Emily Gorman Igor Gorodezky Vincent Guigona Gautam Gupta Nathaniel Hendren Nathan Ilten Sara Kalantari Gary Kilper Marina Kopeliovich Alison Lanski Patrick LaVictoire David Malito Chaitra Nagaraja James Pascaleff Wes Pegden Catherine Pfaff David Purdy Andrew Radin Ryan Reich Danny Rosenthal Anne Shiu Morgan Stern George Sullivan Raymond Tan Benjamin Tradewell Kevin Tucker Jonathan Walsh Kenneth Ward Michael Williams Emily Witt John Zhu

YSP Counsellors

7-8

1. Emily Butler

2. Joyce Chen

3. Emily Gorman

4. Gary Kilper

5. David Malito

6. Chaitra Nagaraja

7. Catherine Pfaff

8. David Purdy

9. George Sullivan

9-10

1. Dean Baskin

2. Gautam Gupta

3. Nathan Ilten

4. Marina Kopeliovich

5. Patrick LaVictoire

6. Wes Pegden

7. Anne Shiu

8. Benjamin Tradewell

9. Emily Witt

11-12

1. Lauren Beitler

2. Matthew Gelvin

3. James Pascaleff

4. Ryan Reich

5. Kevin Tucker

SESAME Counselors

1. Zachary Cohn

2. Nathan Czuba

3. Keith D’Souza

4. Joanna Goldenberg

5. Igor Gorodezky

6. Vincent Guigona

7. Nathaniel Hendren

8. Sara Kalantari

9. Alison Lanski

10. Andrew Radin

11. Danny Rosenthal

12. Morgan Stern

13. Raymond Tan

14. Jonathan Walsh

15. Kenneth Ward

16. Michael Williams

SUMMER 2003 REU PARTICIPANTS AND ASSIGNMENTS 3

17. John Zhu